# Using Mobile AR for Rapid Feasibility Analysis for Deployment of Robots: A Usability Study with Non-Expert Users

### v1

@@ I Introduction
Offline simulation using digital or virtual twins offers a more thorough approach that avoids long downtimes [5]. However, creating such a virtual model incurs high costs, both financial and temporal, as designers need to capture all physical characteristics of the space, recreate them in a virtual environment, and develop complex simulations [6]. While this might be worth the high initial costs for an operational workcell, generating a feasibility analysis might not justify the expense.
@@ II Related work
The deployment of robotic systems involves comprehensive feasibility analysis, which encompasses both technical and economic aspects [10]. Moreover, understanding concepts such as reachability is crucial for optimizing performance and ensuring cost-effectiveness. Here, we review the existing literature on reachability analysis and feasibility studies, highlighting key research that informed our approach. Additionally, we explore methodologies for human-robot interaction (HRI) user studies [11].
@@ II-C Role of User Studies in Human-Robot Interaction
HRI relies on user studies to evaluate the effectiveness and usability of robotic systems. The need for reliable and replicable results in these studies is paramount [11]. Rajendran et al. [28] offer a framework for conducting HRI studies that emphasize replicability and reliability. Fraune et al. [29] provide lessons learned from HRI experts on designing and conducting user studies, stressing the importance of methodological rigour.
User studies are also a common and desirable evaluation approach in AR [30]. They are essential for determining the usability of AR applications, which is crucial for their adoption success [31]. However, many AR user studies are conducted under inadequate experimental conditions and therefore lead to misleading results [32, 33]. Thus, the correct experimental design is crucial for generating high quality results. Researchers utilize user studies to evaluate AR interactions with robotic systems, such as assessing collection interfaces for robot learning [34] or configuring safety zones [35]. For instance, Tadeja et al. [36] conducted user studies with an AR-based system to enhance human-robot collaboration through both formative and quantitative evaluations.
In our approach, we aim to empower expert and non-expert users alike at the pre-deployment phase by helping them make educated decisions about introducing robots into their production lines (see Figure 1). We achieve this by enabling rapid feasibility analysis with minimal costs based on reachability studies employing a context-aware path planner. We acknowledge that more evidence is needed to properly evaluate AR-based interfaces for HRI. Therefore, in this paper, we follow established HRI methodologies for our evaluation and contribute to AR-supported HRI.
@@ III Rapid Feasibility Analysis
[…]
The result of the last step is a successful or not-successful reachability study based on the user input. In case of failure, the user can modify the captured data without the need to restart the entire process.
@@ III-A Tool Implementation
We implemented our prototype of the tool using iPhone 15 Pro Max running iOS 17.6.1 for steps a-e as shown in Figure 2. For the path planning step, we utilized MoveIt2 with the single-query LBKPIECE path planner [38] in ROS2 Indigo. The chosen planner prioritized efficiency and robustness over the optimality of the solution.
@@ IV-A Tasks
The participants were tasked with performing a feasibility analysis of a cobot for two automation tasks using both the proposed tool and a conventional method. The latter involved using RoboDK [39], a widely-used industrial tool known for its simplicity as an automation simulator. The two tasks were:
@@
We have opted not to include end-effectors for either of the tasks. Instead, we used the tool flange as the Tool Center Point in reachability analysis. While the reachability results would be different, adding an end-effector would not increase the complexity of the tasks in terms of the proposed solution.
@@ IV-B Study Protocol
| [min] | ||
| A | Proposed method | 7.24 |
| B | 6.93 | |
| A | RoboDK | 25.81 |
@@ V-B Task Completion Times
As seen in Table I, both tasks were completed on average in 7.09 [min], a significant improvement over RoboDK. Additionally, we set a 30 [min] cutoff time for participants to complete the task using RoboDK, and only two participants managed to do so. To further highlight the difference, the experts who completed the task using RoboDK only managed to model the workcell, skipping the path planning part. All participants attempted to complete task (A) using RoboDK, as it was perceived as the easier task, as indicated in Figure 3.
@@ V-C Prototype Refinement
Improved Pipeline: We fixed ROS2 communication with the to ensure a smooth data handover to MoveIt2.
@@ VII-A Task Completion Times
The average completion times for tasks (A) and (B) were 6.88 [min] and 7.66 [min], respectively, as seen in Table II. For statistical analysis, we used a natural log transformation to normalize the distribution. The Shapiro-Wilk test for normality failed to reject the null hypothesis (), allowing us to use paired t-tests to check for differences between the two tasks. The test revealed no statistical significance between the tasks (). Therefore, we have also used paired Two One-Side t-tests (t-TOST) to test for equivalence. Based on the t-TOST test (), we rejected the presence of effects more extreme than -0.5 to 0.5 [min] (a range chosen based on the desired goal of task completion: 10% of 10 minutes), suggesting that the level of difficulty of the more complex task (B) did not significantly impact the time required to complete it using the proposed method within equivalence interval. Moreover, these times closely align with the experts’ average result of 7.09 [min] (Table I).
@@ VII-C Questionnaires
[…]
As the Shapiro-Wilk test showed deviation from normality (), we employed the Wilcoxon signed-rank test to examine the difference between tasks (A) and (B). The results () indicated no significant difference between the TLX results for tasks (A) and (B), therefore, we have also employed Wilcoxon TOST. The equivalence test with the -10 to 10 range (chosen based on the labels of TLX results, approximately 20 per label) rejected null hypotheses (), indicating that the results are equivalent within the equivalence interval. This shows that task complexity has little to no effect on the cognitive load incurred when using the tool for conducting feasibility studies.
@@ VII-D Observed Behavior and Feedback
Multiple participants faced initial challenges with the bounding box setup, often requiring multiple attempts to set it up correctly. Issues with the bounding box handles were a frequent obstacle, with several participants struggling to manipulate them effectively. These could also lead to slightly decreased but still high average usability scores. Despite these challenges, participants generally enjoyed the tool’s functionalities. For example, they found the 3D scanning and point selection using spraying to be intuitive and satisfying. The visual clarity and responsiveness of the tool were also highlighted positively.
@@ VIII Discussion
The results from the expert and quantitative user studies are strongly positive and indicate a significant improvement in favor of the proposed method over the conventional approaches.
Firstly, the proposed method achieved a substantial time reduction compared to RoboDK. When using the tool, the average completion time was 7.27 [min] for feasibility analysis by non-expert users, whereas RoboDK usage required at least 25.81 [min] by an expert user.
@@
[…]
@@ IX Conclusions
Despite these successes, we see room for improvement. Feedback and observations from participants indicated that certain AR interactions, such as placing a bounding box, require further refinement to enhance intuitiveness. This is reflected in the usability score, i.e., while an average score of is considered very good, some interface refinements could lead to higher ratings. Additionally, while the used tool addresses the reachability aspect of feasibility analysis, other aspects, such as technical and economic feasibility, must also be considered for a comprehensive evaluation.

### v2

@@ I Introduction
Offline simulation using digital or virtual twins offers a more thorough approach that avoids long downtimes [5]. However, creating such models incurs high financial and temporal costs, as designers need to capture all physical characteristics of the space, recreate them virtually, and develop complex simulations [6]. Industrial tools such as RoboDK [7] and Visual Components [8] target expert users in 3D simulation and robotics. While this might justify the high initial costs for an operational workcell, it may not be worth it for a feasibility analysis.
@@ II Related work
The deployment of robotic systems involves comprehensive feasibility analysis, which encompasses both technical and economic aspects [12]. Moreover, understanding concepts such as reachability is crucial for optimizing performance and ensuring cost-effectiveness. Here, we review the existing literature on reachability analysis and feasibility studies, highlighting key research that informed our approach. Additionally, we explore methodologies for human-robot interaction (HRI) user studies from [13].
@@ II-C Role of User Studies in Human-Robot Interaction
HRI relies on user studies to evaluate the effectiveness and usability of robotic systems. The need for reliable and replicable results in these studies is paramount [13]. Rajendran et al. [30] offer a framework for conducting HRI studies that emphasize replicability and reliability. Fraune et al. [31] provide lessons learned from HRI experts on designing and conducting user studies, stressing the importance of methodological rigor.
User studies are also a common and desirable evaluation approach in AR [32]. They are essential for determining the usability of AR applications, which is crucial for their adoption success [33]. However, many AR user studies are conducted under inadequate experimental conditions and, therefore, lead to misleading results [34, 35]. Thus, the correct experimental design is crucial for generating high-quality results. Researchers utilize user studies to evaluate AR interactions with robotic systems, such as assessing collection interfaces for robot learning [36] or configuring safety zones [37]. For instance, Tadeja et al. [38] conducted user studies with an AR-based system to enhance human-robot collaboration through both formative and quantitative evaluations.
In our approach, we aim to empower expert and non-expert users alike at the pre-deployment phase by helping them make educated decisions about introducing robots into their production lines (see Figure 1). We achieve this by enabling rapid feasibility analysis with minimal costs based on reachability studies employing a context-aware path planner. Recognizing the ongoing need for research in AR-supported HRI, this paper evaluates an AR interface for robot reachability assessment using established HRI methodologies, providing insights into the application of how AR can support the crucial initial phase of robot integration within human environments.
@@ III Rapid Feasibility Analysis
Defining Workspace: The user sets the extent of a virtual bounding box to define the environment. The workspace should encompass the automated area, such as for the machine loading task: the opening of the machine, the fixture for the workpiece (like a chuck), and the area where the robot should be positioned. The user can opt to include collision areas, like a wall, in the extended area. Choosing Robot Model and Placement: The user selects the robot size and “places” a virtual robot on a physical or virtual flat surface. An initial reachability sphere associated with a given robot model is displayed to guide users regarding size and placement requirements. Note that this reachability sphere is, at best, a gross approximation of the actual configuration space and does not account for obstacles or other constraints (e.g., singularities), and is only intended as a visual help for the user. 3D Data Collection: The user collects a point cloud representation of the environment within the bounding box. Point Cloud Processing Toolbox: We assume that our solution will be used by users who do not know best practices for 3D scanning, and therefore, our workflow is designed to be iterative and offers a set of tools allowing to modify the captured scan: (i) outlier removal i.e., noise removal from the point cloud; (ii) downsampling, i.e. reduction of points in the point cloud; (iii) brush eraser, i.e., point eraser using a cone-shaped brush on the screen; (iv) sponge eraser, i.e., point eraser using the device as a sponge to “wipe off” points, and (v) object addition, i.e., adding simple primitives to the scan. Interaction Area Markup: The user marks points with a cone-shaped green spray and defines the robot’s approach direction using the device’s orientation relative to the marked area.The user defines as many interaction areas as necessary, such as workpiece source and sink areas and the machine workpiece fixture area, for a machine loading task. Path Planner: The collected data is used to run an online path planner that attempts to reach the specified interaction areas with the robot. The planner employs point cloud data converted into a 3D occupancy map for collision detection and defined robot kinematics to avoid singularities.
The result of the last step is a successful or unsuccessful reachability study based on the user input. In case of failure, the user can modify the captured data without the need to restart the entire process.
@@ III-A Tool Implementation
We implemented our prototype of the tool using iPhone 15 Pro Max running iOS 17.6.1 for steps a-e, as shown in Figure 2. For the path planning step, we utilized MoveIt2 with the single-query LBKPIECE path planner [40] in ROS2 Indigo. The chosen planner prioritized efficiency and robustness over the optimality of the solution.
@@ IV-A Tasks
Participants performed a feasibility analysis of a cobot for two automation tasks using both the proposed tool and RoboDK, a widely-used industrial automation simulator. The tasks were:
@@
We have opted not to include end-effectors in our tasks, as they do not increase user complexity and our focus is on conducting a usability study rather than developing an industry-ready product. However, we plan to add them in the future to account for reachability differences due to the tool flange offset. Adding a gripper will likely limit the reachability region around pick-and-place areas by reducing possible gripping orientations. It will also likely affect the reachability region during arm movement, as the end-effector offset will necessitate different trajectories to avoid collisions.
@@ IV-B Study Protocol
| [min:sec] | ||
| A | Proposed method | 7:14 |
| B | 6:56 | |
| A | RoboDK | 25:49 |
@@ V-B Task Completion Times
As seen in Table I, both tasks were completed on average in 7:05 [min], a significant improvement over RoboDK. Additionally, we set a 30 [min] cutoff time for participants to complete the task using RoboDK, and only two participants managed to do so. To further highlight the difference, the experts who completed the task using RoboDK only managed to model the workcell, skipping the path planning part. All participants attempted to complete task (A) using RoboDK, as it was perceived as the easier task, as indicated in Figure 3.
@@ V-C Prototype Refinement
Improved Pipeline: We improved ROS2 node implementation for smooth data handover to MoveIt2.
@@ VII-A Task Completion Times
The average completion times for tasks (A) and (B) were 6:53 [min] and 7:40 [min], respectively, as seen in Table II. For statistical analysis, we used a natural log transformation to normalize the distribution. The Shapiro-Wilk test for normality failed to reject the null hypothesis (), allowing us to use paired t-tests to check for differences between the two tasks. The test revealed no statistical significance between the tasks (). Therefore, we have also used paired Two One-Side t-tests (t-TOST) to test for equivalence. Based on the t-TOST test (), we rejected the presence of effects more extreme than -30 to 30 [sec] (a range chosen based on the desired goal of task completion: 10% of 10 minutes), suggesting that the level of difficulty of the more complex task (B) did not significantly impact the time required to complete it using the proposed method within equivalence interval. Moreover, these times closely align with the experts’ average result of 7:05 [min] (Table I).
@@ VII-C Questionnaires
The cumulative scores in the TLX questionnaires are shown in Figure 5. The weighted average for task (A) was 30.2, and 31.72 for task (B), which can be both considered “somewhat high”, slightly crossing the bottom threshold of 29 (Table II) [44].
The Shapiro-Wilk test showed deviation from normality (), so we used the Wilcoxon signed-rank test to compare tasks (A) and (B). The results () indicated no significant difference in TLX results between the tasks. We also employed Wilcoxon TOST with equivalence region of -10 to 10, that rejected null hypotheses (), indicating equivalence within the interval. This shows that task complexity has little to no effect on cognitive load when using the tool for feasibility studies.
@@ VII-D Observed Behavior and Feedback
Multiple participants faced initial challenges with the bounding box setup, often requiring multiple attempts to set it up correctly. Issues with the bounding box handles were a frequent obstacle, with several participants struggling to manipulate them effectively. These could also lead to slightly decreased but still high average usability scores. Despite these challenges, participants enjoyed the tool’s functionalities, such as 3D scanning and point selection using spraying, and praised its visual clarity and responsiveness.
@@ VIII Discussion
The expert and quantitative user studies show strong positive results, indicating significant improvement with the proposed method over conventional approaches.
Firstly, the proposed method achieved a substantial time reduction compared to RoboDK. When using the tool, the average completion time was 7:16 [min] for feasibility analysis by non-expert users, whereas RoboDK usage required at least 25:49 [min] by an expert user.
@@
On the negative side, we observed that users struggled with depth perception on a mobile device, indicating a need for simplifying some interface components, such as the bounding box setup. Some users also struggled to scan all important parts of the cell. For example, in task (B), multiple participants failed to scan the top and back of the machine, resulting in failed feasibility analyses because the path planner found trajectories “through” the top of the machine due to the lack of a virtual representation of that part. Conversely, some users collected too many points within the machine and did not clean up the inside, leading to cases where the inside of the machine became unreachable due to floating points, resulting in no viable planning solution. Moreover, using non-iOS-powered mobile devices could lead to broader adoption of our approach but potentially introduce new issues.
@@ IX Conclusions
Despite these successes, we see room for improvement. Feedback and observations from participants indicated that certain AR interactions, such as placing a bounding box, require further refinement to enhance intuitiveness. This is reflected in the usability score, i.e., while an average score of 75.16 is considered very good, some interface refinements could lead to higher ratings. Additionally, while the used tool addresses the reachability aspect of feasibility analysis, other aspects, such as technical and economic feasibility, must also be considered for a comprehensive evaluation.